A student is considering opening a business selling ice cream next summer. The student views this as an alternative to taking summer employment with a local firm where he would earn £4,000 during the 3-month summer period. It would cost £2,000 to obtain a license to operate their stand, £1,200 per month to rent the stand with the necessary equipment and £150 per month for insurance. Petrol costs are estimated at £12 per day and are not affected by sales. The ice cream ingredients can be bought for £5.60 per kilo and cones cost £1.80 per dozen. The ice cream cones contain 100 grams of ice cream and would sell for £1.50 each.

d) Chocolate flakes can be bought at £5 for a box of 20, and the ice creams with flakes are sold for £2.20. If 40% of the customers buy these, calculate the effect of this on the output necessary to make normal profit.
e) If the student can sell 200 ice creams a day, 40% with flakes, estimate the economic profit he would make. Advise him whether he should enter the business.
1
Expert's answer
2020-04-20T18:54:19-0400


1.Calculate fixed costs (TFC):

"2000+1200\times3+150\times3+12\times92=7154"

2.Calculate the variables costs per unit of output (AVC):

flake-free ice cream 100g:

"\frac{5.60}{10}=0.56"

"\frac{1.80}{12}=0.15"

0.56+0.15=0.71

flake ice cream:

"frac{5}{20}=0.25"

0.25+0.71=0.96

e)calculate the breakeven point in pieces

"BEP=\frac{TFC}{P-AVC}""BEP=\frac{7154}{(1.50-0.71)+(2.20-0.96)}=3524"

"0.4\times3524=1409.6=1410"

d) profit:

"200\times92=18400"

"0.4\times18400=7360" flake ice cream

18400-7360=11040 flake-free ice cream

"(11040\times1.50+7360\times2.20)-(11040\times0.71+7360\times0.96)-7154=(16560+16192)-(7838.4+7065.6)-7154=32752-14904-7154=10694"


10694-4000=6694


profit is more than income from employment,

so can do business